### v1

# Subtleties in the pseudomodes formalism

Abstract
The pseudomode method for open quantum systems, also known as the mesoscopic leads approach, consists in replacing a structured environment by a set of auxiliary “pseudomodes” subject to local damping that approximate the environment’s spectral density. Determining what parameters and geometry to use for the auxiliary modes, however, is non-trivial and involves many subtleties. In this paper we revisit this problem of pseudomode design and investigate some of these subtleties. In particular, we examine the scenario in which pseudomodes couple to each other, resulting in an effective spectral density that is no longer a sum of Lorentzians. We show that non-diagonalizability of the pseudomodes’ effective single-particle non-Hermitian Hamiltonian can lead to terms in the effective spectral density which cannot be obtained by diagonalizable non-Hermitian Hamiltonians. We also present a method for constructing the pseudomode parameters to exactly match a fit to a spectral density, and in doing so illuminate the enormous freedom in this process. The case of many uncoupled pseudomodes is explored, and we show how, contrary to conventional assumption, the effective spectral density does not necessarily converge in the limit of an infinite number of pseudomodes; we attribute this to the non-completeness of Lorentzians as basis functions. Finally, we discuss how the notion of effective spectral densities can also emerge in the context of scattering theory for non-interacting systems.
I Introduction
No quantum system can ever be perfectly isolated from its environment, and so understanding the dynamics of open quantum systems is crucial across many fields. Many such systems are coupled sufficiently weakly to their environments that those environments can be approximated by simple Markovian Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equations Gorini et al. (1976); Lindblad (1976). However, many systems of interest couple strongly to and can build up quantum coherences with their environments, from quantum thermal machines Gelbwaser-Klimovsky and Aspuru-Guzik (2015); Thomas et al. (2018); Camati et al. (2020); Ptaszyński (2022); Latune et al. (2023) and solid state devices van der Wiel et al. (2000); Fritz and Vojta (2005); Kim et al. (2022); Ye et al. (2024) to light-matter systems both artificial Anappara et al. (2009); Niemczyk et al. (2010); Agarwal et al. (2013); Forn-Díaz et al. (2019); Qin et al. (2024); Das et al. (2025) and biological Ishizaki and Fleming (2012); Romero et al. (2014); Chen et al. (2015); Scholes et al. (2017). In all these cases, the evolution of the reduced system density matrix becomes non-Markovian, and more advanced techniques are required Tupkary et al. (2022).
[…]
The method has been used to treat problems in areas including fermionic transport Chen et al. (2014); Rams and Zwolak (2020); Wójtowicz et al. (2021); Elenewski et al. (2021); Brenes et al. (2023); Cirio et al. (2023), quantum thermodynamics Lacerda et al. (2023, 2024); Bettmann et al. (2025); Menczel et al. (2024); Albarelli et al. (2024), and quantum optics Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); Lednev et al. (2024); Liang et al. (2024). Recent works have combined this method with tensor networks Brenes et al. (2020); Cirio et al. (2023); Lacerda et al. (2024) and the Time Evolving Density operator with Orthogonal Polynomials Algorithm (TEDOPA) Nüßeler et al. (2022); Ferracin et al. (2024) to treat interacting systems. One advantage of this method is that by modeling the environments using physical modes to form an extended system, we have access to details about correlations between the system and its environment at different energies. This has been used to gain insight into energy-resolved currents for interacting fermionic systems Brenes et al. (2020). Likewise, using the framework of continuous measurement, one can monitor quantum jumps between the residual baths and pseudomodes and perform trajectory unraveling to probe stochastic thermodynamic quantities Bettmann et al. (2025). The auxiliary modes can therefore possess physical significance, rather than serving merely as a mathematical construct for modeling the environment. Alternatively, one can relax the constraint that the pseudomodes have physical parameters and include non-Hermitian terms in the pseudomode Hamiltonian or negative damping rates in the GKSL master equation in order to model an even wider variety of environments Lambert et al. (2019); Luo et al. (2023); Menczel et al. (2024).
The success of the pseudomodes approach, however, relies on the careful design of their geometry and parameters so as to properly mimic the real environment. Implementing this entails various subtleties, which are often overlooked. The goal of this paper is to address said subtleties. In section II, we consider the effective spectral density that arises from a particular configuration of pseudomodes. Depending on the non-Hermitian Hamiltonian generated by the pseudomodes’ Hamiltonian and the residual baths, three possible cases emerge. If none of the pseudomodes couple to each other, then one obtains an effective spectral density that is a sum of Lorentzians, while if the pseudomodes couple to each other there are additional terms in the effective spectral density which we call “Anti-Lorentzian”. We take this one step further, and show that one can obtain terms which are neither Lorentzian nor Anti-Lorentzian by using a non-diagonalizable non-Hermitian Hamiltonian for the pseudomodes. In section III we discuss the reverse process of choosing pseudomode parameters to obtain a particular effective spectral density, which is required to actually implement the pseudomode method. This is generally done in two steps. First, the true spectral density is fit in Fourier space with a set of decaying complex exponentials. Then, the pseudomode parameters must be chosen to match that exponential fit, which is typically done by imposing some additional structure on the pseudomode parameters and then performing optimization Mascherpa et al. (2020). This second step is the greatest challenge to implementing the pseudomode method, in large part because there are many more free parameters for the pseudomodes than there are constraints given by the effective spectral density. We present a new method for choosing the pseudomode parameters, by which the parameters can be chosen to exactly match a decaying exponential fit. Our method also sheds light on the vast number of free parameters involved when choosing the pseudomode parameters. In section IV, we briefly explore the use of uncoupled pseudomodes uniformly distributed in energy and show that, contrary to conventional assumption, the effective spectral density obtained from this approach approximates but does not converge to the true bath spectral density. Instead, we present an alternate method that, while also not convergent in the infinite limit, significantly reduces the error between the true and effective spectral densities in the limit of many pseudomodes and is equally easy to implement. Finally, in section V we show how the effective spectral densities of section II also emerge in the context of scattering theory for non-interacting systems Blanter and Büttiker (2000); Büttiker (1992); Levitov et al. (1996); Esposito et al. (2009). We confirm that if the pseudomode parameters are chosen such that the effective spectral density exactly matches the true spectral density then the currents obtained from Landauer-Büttiker theory will also match exactly between the original system-bath description and pseudomode description. This also presents a new way to study scattering theory currents by breaking down a current between two real baths into the sum of currents between residual baths under a pseudomode mapping.
I.1 Model System
Figure 1: (a) Standard setup of a quantum system coupled to two reservoirs. (b) The pseudomodes model with uncoupled pseudomodes (the Diagonal case), where each bath has been replaced by a finite set of damped pseudomodes. (c) The general pseudomode model, where the pseudomodes are allowed to couple to each other arbitrarily in addition to the residual baths.
We consider the system depicted in Fig. 1(a), consisting of a fermionic system coupled to multiple reservoirs. For simplicity, we take two reservoirs $L,R$ (left and right), but the result generalizes for more than two. The system is described by $n_{S}$ fermionic modes $c_{i}$. The total Hamiltonian of the system and environment is:
$H_{\rm tot}=$ $H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k}\omega_{k}d_{\alpha k}^{\dagger}d_{\alpha k}$ (1)
$+\sum_{\alpha\in\{L,R\}}\sum_{i,k}g_{\alpha i,k}\big(c_{i}^{\dagger}d_{\alpha k}+d_{\alpha k}^{\dagger}c_{i}\big)\,$,
where $d_{\alpha k}$ are the fermionic modes of bath $\alpha\in\{L,R\}$. We set $\hbar=1$ here and throughout the paper. The system Hamiltonian can be completely general and may include arbitrary driving. The last term describes how bath $\alpha\in\{L,R\}$ can, in principle, couple to any mode $c_{i}$ of the system. Each bath is characterized by a spectral density matrix
$(J_{\alpha})_{ij}=2\pi\sum_{k}g_{\alpha i,k}g_{\alpha j,k}\delta(\omega-\omega_{k})\,$, (2)
which is a real and symmetric matrix of dimension $n_{S}\times n_{S}$.
The equation of motion for $c_{i}(t)$, in the Heisenberg picture, reads
$\frac{dc_{i}(t)}{dt}=i[H_{S}(t),c_{i}(t)]-i\sum_{\alpha k}g_{\alpha i,k}e^{-i\omega_{\alpha k}t}d_{\alpha k}$ (3)
$-\sum_{\alpha kj}\intop_{0}^{t}dt^{\prime}g_{\alpha i,k}g_{\alpha j,k}e^{-i\omega_{\alpha k}(t-t^{\prime})}c_{j}(t^{\prime})\,$,
which we can express as a quantum Langevin equation, in terms of noise operators $\xi_{\alpha i}$ and memory kernel matrices $\chi_{\alpha,ij}$, as Brenes et al. (2020):
$\frac{dc_{i}(t)}{dt}=i[H_{S}(t),c_{i}(t)]+\sum_{\alpha}\xi_{\alpha i}(t)-\sum_{\alpha j}\intop_{0}^{t}dt^{\prime}\chi_{\alpha,ij}(t-t^{\prime})c_{j}(t^{\prime})$ (4)
where $\xi_{\alpha i}=-i\sum_{k}g_{\alpha i,k}e^{-i\omega_{\alpha k}t}d_{\alpha k}$ is the noise operator, and the memory kernel is given by the Fourier transform of the spectral density:
$\chi_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}J_{\alpha,ij}(\omega)e^{-i\omega t}\,.$ (5)
While only the $t>0$ part of the memory kernel enters into the dynamics governed by (4), we define the memory kernel for both positive and negative time using (5) so that we can easily convert back and forth between $J_{\alpha,ij}$ and $\chi_{\alpha,ij}$.
The open system dynamics described by (4) are fully determined by the operators $\xi_{\alpha i}(t)$ and $\chi_{\alpha,ij}(t)$. Alternately, one can think of the dynamics as determined by the bath correlation functions:
$C^{+}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)f_{\alpha}(\omega)$ (6)
$C^{-}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)\big(1-f_{\alpha}(\omega)\big)\,$, (7)
where $f_{\alpha}(\omega)$ is the Fermi function associated to bath $\alpha$. From these two expressions we can then see that the dynamics can also be thought of as fully determined by the spectral density matrices $J_{\alpha,ij}(\omega)$ (2) and the Fermi functions $f_{\alpha}(\omega)$ of the baths. This is the most useful description for the implementation of the pseudomode method.
Finally, it is worth noting that if we express the system-bath interaction Hamiltonian as:
$\sum_{\alpha}\sum_{i}c_{i}^{\dagger}B_{\alpha i}+B_{\alpha i}^{\dagger}c\,$, (8)
with $B_{\alpha i}=\sum_{k}g_{\alpha i,k}d_{\alpha k}$ ,then the two correlation functions can equivalently be written as:
$C^{+}_{\alpha ij}(t)=\left\langle B_{i}^{\dagger}(t)B_{j}(0)\right\rangle\,$, (9)
$C^{-}_{\alpha ij}(t)=\left\langle B_{i}(t)B_{j}^{\dagger}(0)\right\rangle\,.$ (10)
I.2 Pseudomode Mapping
Under the pseudomode method, each bath $\alpha\in\{L,R\}$ is replaced by a finite number $n_{\alpha}$ of fermionic pseudomodes $a_{\alpha k}$ (sometimes called lead modes or auxiliary modes), each of which is coupled to its own residual bath, as depicted in Fig. 1(c). The Hamiltonian in (1) is mapped to:
$H_{\rm tot}=H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k,q}\Lambda_{\alpha,kq}a_{\alpha k}^{\dagger}a_{\alpha q}$
$\qquad+\sum_{\alpha\in\{L,R\}}\sum_{i,k}\big(\zeta_{\alpha i,k}a_{\alpha k}^{\dagger}c_{i}+\zeta_{\alpha i,k}^{*}c_{i}^{\dagger}a_{\alpha k}\big)$ (11)
$+\sum_{\alpha\in\{L,R\}}\sum_{k}\sum_{p}\Big\{\omega_{p}b_{\alpha k,p}^{\dagger}b_{\alpha k,p}+\tau_{\alpha k,p}\big(a_{\alpha k}^{\dagger}b_{\alpha k,p}+b_{\alpha k,p}a_{\alpha k}\big)\Big\}\,.$
The first line of (I.2) describes the possible interactions between pseudomodes of the same bath, described by the $n_{\alpha}\times n_{\alpha}$ matrix $\Lambda_{\alpha}$ for each bath. If $\Lambda$ is diagonal then none of the pseudomodes couple to each other, and we have the simpler configuration shown in Fig. 1(b). The second line describes how each pseudomode interacts with the central system through a set of $n_{\alpha}\times n_{S}$ matrices $\zeta_{\alpha}$. Finally, the last line describes a set of infinite dimensional residual reservoirs, one for each pseudomode $a_{\alpha k}$, composed of a set of modes $b_{\alpha k,p}$. The residual baths then have spectral densities given by
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p}\left|\tau_{\alpha k,p}\right|^{2}\delta(\omega-\omega_{p})\,$, (12)
which is a scalar, rather than a matrix like (2), since each residual bath $(\alpha,k)$ couples only to the pseudomode $a_{\alpha k}$. Generally, we want the couplings to the residual baths $\tau_{\alpha k,p}$ to be small compared to the extended system energy levels and constant at all energies, so that the residual baths have constant spectral densities
$J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha k}\,.$ (13)
In this case, the system feels an effective memory kernel given by Mascherpa et al. (2020):
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\zeta_{i}^{\dagger}e^{Wt}\zeta_{j}&t\geq 0\\ \zeta_{j}^{\dagger}e^{-W^{\dagger}t}\zeta_{i}&t\leq 0\,.\end{cases}$ (14)
where
$W_{\alpha}=-i\Lambda_{\alpha}-\Gamma_{\alpha}/2\,.$ (15)
Once again, only the $t\geq 0$ part of $\chi_{\alpha,ij}^{\rm eff}$ enters into the dynamics, but we define the $t<0$ part by $\chi_{\alpha,ij}(-t)=\chi_{\alpha,ij}(t)^{*}$ based on  (5), so that we can take the Fourier transform and conclude that the system feels an effective spectral density given by:
$J^{\rm eff}_{\alpha,ij}(\omega)=2\mathfrak{Im}\left(\zeta^{\dagger}_{\alpha i}\frac{1}{iW_{\alpha}-\omega}\zeta_{\alpha j}\right)\,.$ (16)
The pseudomode parameters are then chosen so that the effective spectral density (16) and effective memory kernel (14) are good approximations for the true spectral density (2) and memory kernel (5) of the original system. The derivation for these forms of $J^{\rm eff}_{\alpha,ij},\chi_{\alpha,ij}^{\rm eff}$ can be found in Appendix A. The reduced system dynamics under (1) will then be identical to those under (I.2) so long as the correlation functions from $J^{\rm eff}_{\alpha,ij}$ are identical to the true correlation functions (6); likewise, the dynamics under (1) will be well-approximated by the dynamics under (I.2) so long as the effective correlation functions well-approximate the true ones Trivedi et al. (2021); Huang et al. (2024). In practice, we achieve this by ensuring that $J^{\rm eff}_{\alpha,ij}(\omega)$ is a good fit to $J_{\alpha,ij}(\omega)$.
The utility of this method thus hinges on our ability to choose a set of $\zeta_{\alpha i}$, $\Lambda_{\alpha}$, and $\Gamma_{\alpha}$ so that the effective spectral density (16) closely matches the actual spectral density (2). In the next section, we describe how to obtain the form of $J^{\rm eff}_{\alpha,ij}$ from the chosen pseudomode configuration; in the following section, we describe the more complicated problem of choosing pseudomode parameters to obtain a desired $J^{\rm eff}_{\alpha,ij}$. As we shall show shortly, there are many possible choices of pseudomode parameters that give rise to the same effective spectral density, and not all of them may be equally desirable. Actually computing the extended-system dynamics generally requires further approximations, most commonly the use of a local GKSL master equation, which may be more or less reasonable to make based on the choice of pseudomode parameters. We do not discuss the merits of particular choices of parameters in this paper, instead focusing on the procedure for constructing the pseudomode parameters and emphasizing the freedoms therein.
II Computation of Effective Spectral Density
The form that the effective spectral density (16) of bath $\alpha$ represented by pseudomodes takes depends on whether the matrix $W_{\alpha}$ (15) is diagonal, diagonalizable, or non-diagonalizable. A pseudomode configuration is said to be diagonal, diagonalizable or non-diagonalizable when its matrix $W_{\alpha}$ has the corresponding property. These cases are described sequentially below. In general we might also have a matrix $W_{\alpha}$ which is block-diagonal, where some blocks are diagonalizable and some are not; in this case, each block can be treated separately and the resulting effective spectral density is merely the sum of the effective spectral densities from each block.
II.1 Diagonal Case
If the pseudomode-pseudomode interaction matrix $\Lambda_{\alpha}$ is diagonal, this means that each pseudomode corresponding to bath $\alpha$ couples to the system but not to any other pseudomodes, as depicted in Fig. 1(b). In this case $e^{W_{\alpha}t}$ is also diagonal. If we label the on-site energies of each pseudomode as $\Lambda_{\alpha,jj}=\varepsilon_{\alpha j}$, and the coupling of each one to their residual baths as $\Gamma_{\alpha,jj}=\gamma_{\alpha j}$, then the resulting effective memory kernel (14) is a sum of complex exponentials:
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\sum_{k}\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}e^{-i\varepsilon_{\alpha k}t-\gamma_{\alpha k}t/2}&t\geq 0\\ \sum_{k}\zeta_{\alpha i,k}\zeta_{\alpha j,k}^{*}e^{-i\varepsilon_{\alpha k}t+\gamma_{\alpha k}t/2}&t\leq 0\,.\end{cases}$ (17)
Taking the Fourier transform, the resulting effective spectral density matrix (16) is then:
$J^{\rm eff}_{\alpha,ij}(\omega)=\sum_{k}\frac{\mathfrak{Re}\left(\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}\right)\gamma_{\alpha k}-2\mathfrak{Im}\left(\zeta_{\alpha i,k}^{*}\zeta_{\alpha j,k}\right)(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (18)
In particular the diagonal elements are:
$J^{\rm eff}_{\alpha,ii}(\omega)=\sum_{k}\frac{\left|\zeta_{\alpha i,k}\right|^{2}\gamma_{\alpha k}}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,$, (19)
which is a sum of Lorentzians with each pseudomode contributing one Lorentzian, with the form:
$\frac{\gamma_{\alpha k}}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (20)
If each bath couples to only one system site, then the effective spectral densities for each bath will contain only a sum of Lorentzians. Any off-diagonal elements contain both Lorentzian terms (20) and terms of the form
$\frac{(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,$, (21)
which we call “Anti-Lorentzian” since they structurally resemble Lorentzians but are anti-symmetric about the energy $\varepsilon$, rather than peaked there.
II.2 Diagonalizable Case
Suppose the matrix $W_{\alpha}$ is not diagonal but can be diagonalized as $W_{\alpha}=S_{\alpha}M_{\alpha}S_{\alpha}^{-1}$ with $M_{\alpha}$ a diagonal matrix. Note that $S_{\alpha}$ is not unitary since $W_{\alpha}$ is not Hermitian. The effective memory kernel matrix (14) is then given by:
$\chi_{\alpha,ij}^{\rm eff}(t)=\zeta_{\alpha i}^{\dagger}S_{\alpha}e^{{M_{\alpha}}t}S_{\alpha}^{-1}\zeta_{\alpha j}\,.$ (22)
Labeling the eigenvalues of $W_{\alpha}$ as $i\varepsilon_{j}-\gamma_{j}/2$ and writing out the multiplication explicitly, we obtain:
$\chi_{\alpha,ij}^{\rm eff}(t)=\begin{cases}\sum_{k}\kappa_{\alpha k,ij}e^{-i\varepsilon_{\alpha k}t-\gamma_{\alpha k}t/2}&t\geq 0\\ \sum_{k}\kappa_{\alpha k,ij}^{*}e^{-i\varepsilon_{\alpha k}t+\gamma_{\alpha k}t/2}&t\leq 0\end{cases}\,$, (23)
with the amplitudes
$\kappa_{\alpha k,ij}=\sum_{pq}\zeta_{\alpha i,p}^{*}\zeta_{\alpha j,q}S_{\alpha,pk}S_{\alpha,kq}^{-1}\,.$ (24)
Note that the diagonal elements $\kappa_{\alpha k,ii}$ always sum to a positive real number:
$\sum_{k}\kappa_{\alpha k,ii}>0\,.$ (25)
Under this model, even if bath $\alpha$ couples to only one system site, it is still possible to obtain $\mathfrak{Im}(\kappa_{\alpha k,ii})\neq 0$. The effective spectral density (16) thus has a similar form to the diagonal case, but because the amplitudes (24) are more complicated, it can now have both Lorentzian (20) and Anti-Lorentzian (21) terms in both its diagonal and off-diagonal elements:
$J^{\rm eff}_{\alpha,ij}(\omega)=\sum_{k}\frac{\mathfrak{Re}\left(\kappa_{\alpha k,ij}\right)\gamma_{\alpha k}-2\mathfrak{Im}\left(\kappa_{\alpha k,ij}\right)(\omega-\varepsilon_{\alpha k})}{(\omega-\varepsilon_{\alpha k})^{2}+(\gamma_{\alpha k}/2)^{2}}\,.$ (26)
Figure 2: (a) Comparison of fits to an effective spectral density from a diagonal configuration (sum of Lorentzians, with uncoupled pseudomodes) and from a diagonalizable one (sum of Lorentzians and Anti-Lorentzians, with pseudomodes coupled to one another). The true spectral density is $J(\omega)=\sqrt{1-\omega^{2}}$, which both fits model using 6 pseudomodes. The diagonal-case fit was obtained by brute-force optimization for the parameters $|\zeta_{k}|^{2},\varepsilon_{k},\gamma_{k}$, while for the diagonalizable case, Prony’s method (described in section III and appendix D) was used to obtain the best-fit $\kappa_{k},\varepsilon_{k},\gamma_{k}$. (b) Contributions to the effective spectral density in the diagonal case from each of the 6 pseudomodes; each mode contributes a Lorentzian term. (c) Contributions to the effective spectral density in the diagonalizable case from each of the 6 pseudomodes; each mode contributes the sum of a Lorentzian term and an Anti-Lorentzian term.
Using a diagonalizable but not diagonal $W$ often allows for a much better fit to structured spectral densities than using a diagonal $W$, as the negativities and anti-symmetry from the Anti-Lorentzian terms allow one to capture features much more easily. Fig. 2 compares 6-mode fits to the semi-elliptical spectral density, showing how the diagonalizable version matches the true spectral density much more closely, particularly around the sharp corners at $\omega=\pm 1$.
II.3 Non-Diagonalizable Case
So long as $W_{\alpha}$ is diagonalizable, we can only end up with Lorentzian (20) and Anti-Lorentzian (21) terms in the spectral density matrix. However, in the case that $W_{\alpha}$ is non-diagonalizable (sometimes called “defective”), other functional forms can appear. An $n\times n$ matrix is non-diagonalizable if it has fewer than $n$ unique eigenvectors, and is “maximally defective” if it has only one eigenvector.
If $W_{\alpha}$ is non-diagonalizable, it can be written in Block Jordan form. Each eigenvalue of $W_{\alpha}$ corresponds to a $d_{i}\times d_{i}$ block, where $d_{i}$ is the algebraic multiplicity of that eigenvalue. For simplicity suppose that $W_{\alpha}$ has one maximally defective eigenvalue $w$; the general result then follows. In that case we can write
$W=S(w\mathbf{1}+N)S^{-1}=w\mathbf{1}+SNS^{-1}\,$, (27)
where $N_{ij}=\delta_{i,j-1}$ is a matrix with all $1$s on the superdiagonal and zeros elsewhere. Then $e^{Wt}=e^{wt}Se^{Nt}S^{-1}$ where the Taylor series for $e^{N}$ only contains $n$ terms since $N$ is nilpotent. We can use this to compute the form of $e^{N}$ as:
$\left(e^{Nt}\right)_{i\geq j}=\frac{t^{i-j}}{(i-j)!}\,.$ (28)
The resulting contribution to the effective memory kernel (14) (for $t>0$) is then:
$\chi_{\alpha,ij}^{\rm eff}(t)=e^{wt}\sum_{p,q}\sum_{k=1}^{n}\sum_{l=1}^{k}\frac{1}{(k-l)!}\zeta_{\alpha i,p}^{*}\zeta_{\alpha j,q}S_{pk}S^{-1}_{lq}t^{k-l}\,.$ (29)
We can see in particular that since $1\leq l\leq k\leq d$, the highest possible power of $t$ that we can get inside the summation is $t^{d-1}$. The general form of the memory kernel is thus
$\chi(t)\sim\sum_{k=0}^{d-1}\kappa_{k}t^{d}e^{wt}\,.$ (30)
[…]
$\frac{1}{(\omega^{2}+\gamma^{2})^{k+1}}\,.$ (31)
Examples of the types of terms that arise are shown in table 1. Note that while there are known methods for constructing $n\times n$} non-diagonalizable matrices Scott (1997), the form of (29) makes it difficult to construct such matrices that will give rise to a memory kernel that is an arbitrary linear combination of such terms.
Table 1: Contribution to the effective spectral density from terms $t^{k}e^{wt}$ in the effective memory kernel for $t>0$. Note that terms with real and imaginary coefficients give different functional forms since $\chi(-t)=-\chi(t)^{*}$}.
Term in $\chi_{\rm eff}(t)$ Term in $J_{\rm eff}(\omega)$
$e^{-\gamma t}$ $\frac{2\gamma}{\omega^{2}+\gamma^{2}}$
$ie^{-\gamma t}$ $\frac{2\omega}{\omega^{2}+\gamma^{2}}$
$te^{-\gamma t}$ $\frac{2(\gamma^{2}-\omega^{2})}{(\omega^{2}+\gamma^{2})^{2}}$
$ite^{-\gamma t}$ $\frac{4\gamma\omega}{(\omega^{2}+\gamma^{2})^{2}}$
$t^{2}e^{-\gamma t}$ $\frac{4(\gamma^{3}-3\gamma\omega^{2})}{(\omega^{2}+\gamma^{2})^{3}}$
$it^{2}e^{-\gamma t}$ $\frac{4(3\gamma^{2}\omega-\omega^{3})}{(\omega^{2}+\gamma^{2})^{3}}$
$t^{n}e^{-\gamma t}$ $n!\left[(\gamma+i\omega)^{-n-1}+(\gamma-i\omega)^{-n-1}\right]$
$it^{n}e^{-\gamma t}$ $in!\left[(\gamma-i\omega)^{-n-1}-(\gamma+i\omega)^{-n-1}\right]$
Example: 2x2 non-diagonalizable Case
Consider the case where we map bath $\alpha$, which couples to only one system site (so the spectral density matrix is 1x1) to two pseudomodes, so $W_{\alpha}$ is 2x2. If we want $W_{\alpha}$ to be non-diagonalizable, the most general form we can pick is:
$W_{\alpha}=\begin{bmatrix}-i\varepsilon-\eta&-i\delta\\ -i\delta&-i\varepsilon-\eta-2\delta\end{bmatrix}\,$, (32)
with system couplings $\zeta=\begin{bmatrix}\zeta_{1}&\zeta_{2}\end{bmatrix}^{T}$}. The effective memory kernel (14) then has an exponential component and a component that is $t$ times an exponential:
$\chi_{\alpha}^{\rm eff}(t)=e^{-i\varepsilon t}e^{-(\delta+\eta)t}\Big\{\left(|\zeta_{1}|^{2}+|\zeta_{2}|^{2}\right)+$ (33)
$\delta t\left[|\zeta_{1}|^{2}-|\zeta_{2}|^{2}-i\left(\zeta_{1}\zeta_{2}^{*}+\zeta_{1}^{*}\zeta_{2}\right)\right]\Big\}\,.$ (34)
The resulting effective spectral density (16) then has three terms: a Lorentzian, an Anti-Lorentzian, and a term that looks like a squared Lorentzian plus a squared Anti-Lorentzian:
$J^{\rm eff}_{\alpha}(\omega)=\left(\left|\zeta_{1}\right|^{2}+\left|\zeta_{2}\right|^{2}\right)\frac{2(\delta+\eta)}{(\omega-\varepsilon)^{2}+(\delta+\eta)^{2}}$
$+(\zeta_{1}\zeta_{2}^{*}+\zeta_{1}^{*}\zeta_{2})\frac{4\delta(\delta+\eta)(\omega-\varepsilon)}{\left[(\omega-\varepsilon)^{2}+(\delta+\eta)^{2}\right]^{2}}$ (35)
$+\left(\left|\zeta_{1}\right|^{2}-\left|\zeta_{2}\right|^{2}\right)\frac{2\delta\left[(\delta+\eta)^{2}-(\omega-\varepsilon)^{2}\right]}{\left[(\omega-\varepsilon)^{2}+(\delta+\eta)^{2}\right]^{2}}\,.$
If we take the case where $\eta=0$ and $\zeta_{2}=0$, then this $W$ represents two pseudomodes in a chain-like configuration, with the system coupled to mode $1$, which is then coupled to mode $2$, which is coupled to a residual bath, as shown in Fig. 3. In this case, the spectral density we obtain is a squared Lorentzian:
$J^{\rm eff}_{\alpha}(\omega)=\frac{4\delta^{3}|\zeta_{1}^{2}|}{\left[(\omega-\varepsilon)^{2}+\delta^{2}\right]^{2}}\,.$ (36)
Figure 3: The simplest non-diagonalizable pseudomode configuration, consisting of two pseudomodes. One mode is coupled to the system and the other is coupled to the residual bath. If the energies and couplings are chosen so that $W$ is non-diagonalizable, then the system feels an effective spectral density which is a squared Lorentzian, as in (36).
We can generalize this to the case where we have several pairs of pseudomodes, where each pair has its own $\varepsilon_{k},\delta_{k},\zeta_{k},\eta_{k}$ and obtain a spectral density that is a sum of terms of the form (II.3).
These non-diagonalizable configurations allow for more variety in the types of effective spectral densities that can be obtained from the pseudomode method beyond only Lorentzians and Anti-Lorentzians, though the terms obtained this way can be closely approximated by just Lorentzians and Anti-Lorentzians. If we treat a non-diagonalizable configuration as the limit of a barely-diagonalizable configuration, then the resulting diagonalizable-type spectral density approaches a non-diagonalizable one, as shown in Appendix C. However, these non-diagonalizable configurations also provide clues into how to fit more general types of spectral densities. For example, from Table 1 we can see that a spectral density whose denominator contains terms of order up to $\omega^{6}$ might be able to be fit exactly by just 3 modes.
III Few-Mode Fitting
While determining the effective spectral density due to a particular pseudomode configuration is fairly straightforward, what we generally want to do is the reverse process: start with a bath with a known spectral density and determine the pseudomode parameters that best model that bath. Specifically, we wish to choose the pseudomode parameters with the following goals:
  1. (i)
    The effective spectral density matrix $J^{\rm eff}_{\alpha,ij}(\omega)$ should closely match the true spectral density matrix $J_{\alpha,ij}(\omega)$.
  2. (ii)
    The couplings to the residual baths should be as small as possible, to justify the use of a local master equation approach to describe them.
  3. (iii)
    The number of pseudomodes used should be as low as possible to make the problem more amenable to computer simulations.
Items (ii) and (iii) exist somewhat in conflict, since if the $J_{\alpha,ij}$ are not sharply peaked then the fewer pseudomodes we use, the larger the couplings between those pseudomodes and the residual baths will be.
Depending on whether we prioritize (ii) or (iii) more highly, two approaches exist. If we prioritize using a small number of modes, as is most common, we first perform a fit to the memory kernel and then try to find pseudomode parameters that give the correct amplitudes. This procedure is discussed in the remainder of this section. If instead we are willing to use a large number of pseudomodes to ensure the residual bath couplings will be small, one option is to proceed similarly to Refs. Brenes et al. (2020); Lacerda et al. (2023, 2024); Bettmann et al. (2025) as described in section IV.
We will assume a diagonalizable configuration for this section where not otherwise specified. Thus, the effective memory kernel is given by (22) and the effective spectral density is given by (26).
III.1 Fitting the Memory Kernel with Prony’s Method
We fit the memory kernel from each bath separately. We first fit each element of $\chi_{\alpha,ij}^{\rm eff}(t)$ as a sum of complex exponentials using Prony’s method Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022), an algorithm originally designed for discrete signal processing and detailed in Appendix D. Prony’s method works by fitting a discrete set of evenly spaced points with complex exponentials; we found the best results by performing optimization of the Prony’s method fit for a variable set of discrete points. This gives us the set of $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$ from which to construct the effective memory kernel (22). Since the exponents in $\chi_{\alpha,ij}^{\rm eff}$ are independent of $i,j$, in the end each $i,j$ element of the memory kernel matrix should be fit by a sum of exponentials with the same exponents but different amplitudes $\kappa_{\alpha k,ij}$. Caution must be used here, since the fit obtained may produce a spectral density that is not positive everywhere or even has non-decaying exponentials. Few-mode fitting has also been performed by brute-force optimization over the pseudomode parameters without first fitting the memory kernel Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); doing so guarantees a fit with physical pseudomode parameters, but is much less efficient computationally. Note, also, that some spectral densities will be easier to fit than others. In particular, if we have a wideband $J(\omega)=\Gamma_{0}$ for $\omega_{m}\leq\omega\leq\omega_{M}$ and 0 otherwise, then the corresponding memory kernel is
$\chi(t)=\frac{i\Gamma_{0}}{t}\left(e^{-i\omega_{M}t}-e^{-i\omega_{m}t}\right)\,.$ (37)
Because this memory kernel decays as $1/t$, it cannot be well-fit by any finite set of decaying exponentials. Thus, somewhat ironically, a large but flat spectral density is a very difficult problem to describe with the pseudomode approach.
III.2 Obtaining Pseudomode Parameters
Choosing the parameters $\Lambda,\Gamma,\zeta$ to obtain the desired amplitudes $\kappa$ is an Inversion Problem, and it remains the greatest challenge in implementing the pseudomode technique. The relation (24) between the constants $\kappa_{\alpha k,ij}$ and the actual pseudomode parameters $\Lambda_{\alpha},\Gamma_{\alpha},\zeta_{\alpha}$ is highly nonlinear, and there are many more degrees of freedom in these parameters than are required. The fit with Prony’s method for $n$ pseudomodes gives us $4n$ real parameters: $n$ complex amplitudes and $n$ complex exponents. For the actual pseudomode parameters, we have $n^{2}+3n$ real free parameters: $n^{2}$ from $\Lambda$ (Hermitian), $n$ from $\Gamma$ (real and diagonal), and $2n$ from $\zeta$ (complex). However, not all of these free parameters actually contribute due to the structure of the $\kappa$’s (24), and there are a number of constraints that are not readily apparent. For example, a simple one is that, because we always see pairs $\zeta_{\alpha i}^{*}\zeta_{\alpha j}$, the vector $\zeta_{\alpha i}$ is invariant under a global phase $\zeta\to e^{i\phi}\zeta$.
We can formally express this Inversion Problem as follows. For this section we will suppress the indices $i,j$ and consider the case where each bath only couples to one system site; we will also suppress the bath index $\alpha$ since each bath is fit independently. Given a fit consisting of a set of $\kappa_{k},\varepsilon_{k},\gamma_{k}$, our goal is to construct $W$ and $\zeta$. Now,
$W=SMS^{-1}\,,\quad M_{ij}=\delta_{ij}(-i\varepsilon_{i}-\gamma_{i}/2)\,$, (38)
where $M$ is a diagonal matrix whose entries are the eigenvalues of $W$ obtained from the Prony’s method fit. While the eigenvalues of $W$ are specified, $S$ is not. Note that if we transform $W$ by some unitary transformation, then as long as we transform $\zeta$ by the same unitary we will obtain the same effective memory kernel (22):
$\chi^{\rm eff}_{ij}(t)$ $=(U\zeta_{i})^{\dagger}USe^{Mt}S^{-1}U^{\dagger}(U\zeta_{j})$ (39)
$=\zeta_{i}^{\dagger}Se^{Mt}S^{-1}\zeta_{j}\,.$
We must then choose the elements of the matrix $S$ and vector $\zeta$ to get the right amplitudes $\kappa$ according to (24).
We can rewrite the equation for the $\kappa_{k}$ in a more useful form as:
$\kappa_{k}$ $=(\zeta^{\dagger}S)_{k}(S^{-1}\zeta)_{k}$ (40)
$=(v^{\dagger})_{k}u_{k}\,$, (41)
where we define the vectors
$u=S^{-1}\zeta\,,\qquad v=S^{\dagger}\zeta.$ (42)
Note that the relation between these two vectors is
$S^{\dagger}Su=v\,.$ (43)
Thus, we can express the Inversion Problem as the problem of, given a vector $\kappa_{k}$, choosing two vectors $u,v$ such that
$v_{k}^{*}u_{k}=\kappa_{k}\,,\quad u^{\dagger}v>0\,$, (44)
which are related by a positive semi-definite matrix $S^{\dagger}S$. Notice that $u^{\dagger}v=\sum_{k}\kappa_{k}>0$ is guaranteed from (25). Equation (44) is now bi-linear in $u,v$. We have thus expressed this as a Bi-linear Inversion Problem (BIP), a class of problems that tend to be highly underdetermined Li et al. (2015); Aghasi et al. (2018).
We can write any solution to (44) by first fixing the values of $u_{j}$ and then taking
$v_{j}=(\kappa_{j}/u_{j})^{*}\,.$ (45)
Our task is now to construct a matrix $S^{\dagger}S$ which is positive definite and satisfies (43). We can see such a matrix for $n$ pseudomodes will have the form
[…]
where $B_{\perp}$ can be any positive matrix that satisfies $B_{\perp}u=0$ and makes (46) Hermitian.
We will use a convenient change of basis to construct a matrix of this form. For $n$ pseudomodes, we choose any vectors $b_{1},...,b_{n-2}$ such that the set $\{v,u,b_{1},...,b_{n-2}\}$ is linearly independent. We then perform Graham-Schmidt orthonormalization on this set to obtain a new orthonormal basis $\{e_{1},e_{2},...,e_{n}\}$. In this basis, note that by construction we have
$v=||v||e_{1}\,,\quad u=(u.e_{1})e_{1}+(u.e_{2})e_{2}\,$, (47)
where $u$ and $v$ only have one or two nonzero elements. The unitary matrix that implements this change of basis is the matrix whose columns are the $e_{j}$:
$U_{e}=\begin{bmatrix}|&|&&|\\ e_{1}&e_{2}&\dots&e_{n}\\ |&|&&|\end{bmatrix}\,$. (48)
Thus, we can construct the matrix $S^{\dagger}S$ as block-diagonal in the $e_{j}$ basis with two blocks
$S^{\dagger}S=U_{e}\begin{bmatrix}A&\mathbf{0}\\ \mathbf{0}&B\end{bmatrix}U_{e}^{\dagger}\,$, (49)
where the block $A$ is $2\times 2$ and satisfies
[…]
$A_{12}^{*}u_{1}+A_{22}u_{2}=0\,$,
and the block $B$ is $(n-2)\times(n-2)$ and positive definite. This is analogous to the decomposition in (46). A simple choice is to take $B=\mathds{1}_{n-2}$ as the identity matrix. The matrix $A$ is then constructed by solving (LABEL:eqn:A_block) for $A_{11},A_{22}$; we then choose $A_{12}$ so that $S^{\dagger}S$ is positive semi-definite, which requires
$|A_{12}|^{2}<A_{11}A_{22}\,.$ (51)
Once we have constructed the blocks $A,B$, we can then obtain $S^{\dagger}S$ from (49). Since $S^{\dagger}S>0$, its square root $S=\sqrt{S^{\dagger}S}$ exists, is well-defined, and is Hermitian. From there, we use (42) and (38) to construct
[…]
In general, the matrix $\tilde{\Gamma}$ will not be diagonal at this stage. Thus, we define $U_{\Gamma}$ as the unitary matrix that diagonalizes $\tilde{\Gamma}$
$\Gamma=U_{\Gamma}\tilde{\Gamma}U_{\Gamma}^{\dagger}\,$, (53)
and then rotate both $\zeta$ and $W$ by this same unitary:
$W=U_{\Gamma}\tilde{W}U_{\Gamma}^{\dagger}=U_{\Gamma}SMS^{-1}U_{\Gamma}^{\dagger}\,$, (54)
$\zeta=U_{\Gamma}\tilde{\zeta}=U_{\Gamma}Su\,.$ (55)
We can now finally construct the actual pseudomode parameters $\Lambda,\Gamma,\zeta$ corresponding to a set of $\kappa_{k},\varepsilon_{k},\gamma_{k}$ obtained from Prony’s method:
$\Lambda=\frac{1}{2i}(W^{\dagger}-W)\,,\quad\Gamma=-(W+W^{\dagger})\,.$ (56)
This approach can always be used to obtain a set of pseudomode parameters whose effective spectral density will exactly match the effective spectral density obtained from fitting using Prony’s method, provided that Prony’s method outputs a set of $\kappa_{k}$ that satisfy (25). However, this approach is not guaranteed to give a matrix $\Gamma>0$.
It is worth emphasizing again here the sheer amount of freedom involved in this process. In (45) we can choose anything for the $u_{j}$. Then in (46), $B_{\perp}$ can by any positive-definite $n\times n$ matrix satisfying $B_{\perp}u=0$; alternately, we can choose any positive block $B$ in (49) and any value of $A_{12}$ satisfying (51). A naturally interesting question is then: for what fit parameters $\kappa_{k},\varepsilon_{k},\gamma_{k}$ will there be a choice of $u_{j},B_{\perp}$ such that we obtain $\Gamma>0$? We were unable to find a general answer to this question, but we explore it further for the simplest case of 2 pseudomodes in Appendix B. For that case, we show that it is possible to have $\Gamma>0$ if and only if the effective spectral density from the fit parameters is positive everywhere.
IV Many-Mode Fitting
If we are willing to use a large number of pseudomodes, we can ensure that couplings to the residual baths remain small. However, it stops being practical to run optimization algorithms for fitting. In this case, we might proceed as in Refs. Brenes et al. (2020); Lacerda et al. (2023, 2024); Bettmann et al. (2025). The procedure uses a diagonal configuration, so each pseudomode couples only to the system and to its own residual bath and the effective spectral density is given by (18). There are a few benefits to this many-mode procedure. It is more efficient to set up since no optimization routine is required, though computation with a large number of modes is expensive. Additionally, by using a large number of uncoupled modes at different energies, analyzing the currents between individual pseudomodes and the system can provide information about energy currents. For each bath, the procedure is as follows, where the index $\alpha$ is suppressed:
  1. 1.
    Define a relevant frequency range $\omega_{m},\omega_{M}]$ over which to approximate the spectral density.
  2. 2.
    Choose a number of pseudomodes $n$ such that
    $\gamma=\frac{\omega_{M}-\omega_{m}}{d-1}$ (57)
    is an acceptably small coupling to the residual baths.
  3. 3.
    Let $\varepsilon_{j}=\omega_{m}+(j-1)\gamma$, such that the $\varepsilon_{j}$ form an evenly spaced array with $\varepsilon_{1}=\omega_{m}$ and $\varepsilon_{d}=\omega_{M}$.
  4. 4.
    Construct $\Lambda,\Gamma$ by $\Lambda_{ij}=\delta_{ij}\varepsilon_{j}$ and $\Gamma_{ij}=\delta_{ij}\gamma$.
  5. 5.
    Choose the elements of $\zeta$ so that
    $\zeta_{i,k}\zeta_{j,k}=J_{ij}(\varepsilon_{k})\gamma/2\pi$.
Since the matrices $\Lambda,\Gamma$ are diagonal, the effective spectral density matrix is given by (18). Furthermore, since all the $\zeta_{ik}$ are real, the resulting elements of the spectral density matrix $J_{ij}$ will be sums of Lorentzians centered at the elements $\varepsilon_{k}$ with width $\gamma/2$:
$J^{\rm eff}_{ij}(\omega)=\sum_{k=1}^{n}\frac{J_{ij}(\varepsilon_{k})}{2\pi}\frac{\gamma^{2}}{\left(\omega-\varepsilon_{k}\right)^{2}+(\gamma/2)^{2}}\,.$ (58)
It has been claimed that this sum will converge to $J_{\alpha,ij}(\omega)$ as $n\to\infty$ (assuming we adjust $\gamma$ accordingly); however, this is not the case. Directly applying the procedure described above for a large number of modes to a semi-elliptical spectral density yields the effective spectral density depicted in Fig. 4, which oscillates rapidly about the true spectral density.
Figure 4: Effective spectral density obtained by applying the many-mode approximation procedure described above to a semi-elliptical spectral density, for 100 pseudomodes. The effective spectral density oscillates rapidly about the true spectral density.
IV.1 Proof of Nonconvergence
The problem of nonconvergence arises because the spacing between Lorentzians is also used as the width of the Lorentzians, and so the ratio between their widths and separations remains constant rather than tending to zero and we cannot reduce the Lorentzians to delta functions.
To show where the issue emerges, consider $J^{\rm eff}_{ij}(\varepsilon_{k})$, the effective spectral density at one of the pseudomode energies. First, note that we can write:
$J^{\rm eff}_{ij}(\varepsilon_{k})=\sum_{q=1}^{n}\frac{J_{ij}(\varepsilon_{q})}{2\pi\left[(q-k)^{2}+1/4\right]}\,$, (59)
where the $\gamma$’s conveniently cancel since $\varepsilon_{q}-\varepsilon_{k}=(q-k)\gamma_{\alpha}$. Now, assuming a very large number of pseudomodes, we can approximate $J^{\rm eff}_{ij}(\varepsilon_{q})$ in (59) by
$J^{\rm eff}_{ij}(\varepsilon_{q})\approx J^{\rm eff}_{ij}(\varepsilon_{k})+J^{\prime}(\varepsilon_{k})(q-k)\gamma\,$, (60)
where the shorthand $J^{\prime}=dJ/d\omega$. We then can see that the contributions from the linear terms at $\varepsilon_{k+q}$ and $\varepsilon_{k-q}$ will cancel each other, so we can rewrite (59) as:
$J^{\rm eff}_{ij}(\varepsilon_{k})=\sum_{q}\frac{J_{ij}(\varepsilon_{k})}{2\pi\left[(q-k)^{2}+1/4\right]}\,.$ (61)
We can arrive at (61) more rigorously if we restrict the sum to a set number of terms ($q$ such that $|k-q|\leq 8$ is sufficient to show nonconvergence) and use the fact that $J_{ij}$ is a continuous function to show that we can always choose a number of modes so that $J_{ij}(\varepsilon_{k})-J_{ij}(\varepsilon_{q})<\epsilon$ for any $\epsilon$. In the limit of infinitely many pseudomodes we can take $\epsilon\to 0$ and the limits on the summation to $\pm\infty$ and likewise arrive at (61).
In the limit where the number of pseudomodes goes to infinity, we can then reindex the summation in (61) as:
$J^{\rm eff}_{ij}(\varepsilon_{k})=\frac{J_{ij}(\varepsilon_{k})}{2\pi}\sum_{\ell=-\infty}^{\infty}\frac{1}{\ell^{2}+1/4}\,$, (62)
and finally evaluating the summation:
$J^{\rm eff}_{ij}(\varepsilon_{k})=J_{ij}(\varepsilon_{k})\coth(\pi/2)\approx 1.09J_{ij}(\varepsilon_{k})\,.$ (63)
We can use this same process to find the complete effective spectral density in the infinite-mode limit with this same process. If we let $r$ be between 0 and $1$ then we can write $J^{\rm eff}_{ij}(\varepsilon_{j}+r\gamma)=\eta_{1}(r)J_{ij}(\varepsilon_{j}+r\gamma)$ where using the same technique to reach an infinite summation we now have:
$\eta_{1}(r)$ $=\frac{1}{2\pi}\left({\sum_{\ell=0}^{\infty}\frac{1}{(\ell+r)^{2}+\frac{1}{4}}}+\sum_{\ell=1}^{\infty}\frac{1}{(\ell-r)^{2}+\frac{1}{4}}\right)$ (64)
$=\frac{-\sinh(\pi)}{\cos(2\pi r)-\cosh(\pi)}\,$, (65)
which we can see is periodic in $r$ with period 1. We can see these oscillations in Fig. 5.
Figure 5: Effective spectral density for pseudomodes in a diagonal configuration (blue) and for modes in 2x2 non-diagonalizable blocks (pink), compared to the error expected in the infinite-mode limit at the pseudomode energies (above $J_{\rm eff}$) and halfway between them (below $J_{\rm eff}$, for two example spectral densities. 100 pseudomodes are used in each figure, creating an effective spectral density that is either the sum of 100 Lorentzians or 50 squared Lorentzians. Note that where the effective spectral density differs from the infinite-mode prediction, it is where the second derivative of the true spectral density is large, so the first-order Taylor approximation (60) is locally a poor approximation. The left spectral density is semi-elliptical, and the right is a sum of two Lorentzians.
IV.2 Many-Mode Approach with Non-diagonalizability
We can do slightly better in the large $n$ limit at the cost of doubling the number of modes by replacing the pseudomodes above with 2x2 non-diagonalizable blocks, with one of the two pseudomodes in each block coupled to the system and the other coupled to the residual bath as shown in Fig. 3, such that each block contributes a squared Lorentzian to the effective spectral density as in (36). The steps above are modified as:
  1. 1.
    Define a relevant frequency range $\omega_{m},\omega_{M}]$ over which to approximate the spectral density
  2. 2.
    Choose an even number of pseudomodes $n$ such that $\gamma=4\left(\frac{\omega_{M}-\omega_{m}}{d/2-1}\right)$ is an acceptably small coupling to the residual baths.
  3. 3.
    Let $\delta=\gamma/4$ and $\varepsilon_{q}=\left[\omega_{m}+\delta(q-1)\right]$ for q=1,,n/2{q=1,...,n/2}, such that the $\varepsilon_{q}$ form an evenly spaced array with $\varepsilon_{1}=\omega_{m}$} and $\varepsilon_{d/2}=\omega_{M}$}.
  4. 4.
    Construct $\Lambda$ and $\Gamma$ as block-diagonal, with $n/2$ 2x2 blocks $\Lambda^{q}$ and $\Gamma^{q}$ of the form:
    $\Lambda^{q}=\begin{bmatrix}\varepsilon_{q}&\delta\\ \delta&\varepsilon_{q}\end{bmatrix}\,,\qquad\quad\Gamma^{q}=\begin{bmatrix}0&0\\ 0&4\delta\end{bmatrix}\,$. (66)
  5. 5.
    Choose the elements of $\zeta$ so that
    $\zeta_{i,k}\zeta_{j,k}=J_{ij}(\varepsilon_{(k+1)/2})\gamma/2\pi$ if $k$ is odd and 0 if $k$ is even.
Under this method, we obtain an effective spectral density that is a sum of $n/2$ squared Lorentzians:
$J^{\rm eff}_{ij}(\omega)=\sum_{k=1}^{n/2}\frac{4J_{ij}(\varepsilon_{k})}{2\pi}\frac{\delta^{4}}{\left[(\omega-\varepsilon_{k})^{2}+\delta^{2}\right]^{2}}\,.$ (67)
There is still error in the large $n$ limit for this method, but it is smaller than that for the one-layer version. If we proceed analogously to before, we obtain at the pseudomode energies
$J^{\rm eff}_{ij}(\varepsilon_{k})$ $=\frac{2J_{ij}(\varepsilon_{k})}{\pi}\sum_{\ell=-\infty}^{\infty}\frac{1}{(\ell^{2}+1)^{2}}$ (68)
$=\frac{2}{\pi}\big[\coth(\pi)+\pi\mathrm{csch}(\pi)\big]J_{\alpha,ij}(\varepsilon_{k})$ (69)
$\approx 1.027J_{ij}(\varepsilon_{k})\,$, (70)
and in general we obtain $J^{\rm eff}_{ij}(\varepsilon_{k}+r\gamma)=\eta_{2}(r)J_{ij}(\varepsilon_{k}+r\gamma)$ with
$\eta_{2}(r)$ $=\frac{2}{\pi}{\left(\sum_{\ell=0}^{\infty}\frac{1}{\left[(\ell+r)^{2}+1\right]^{2}}+\sum_{\ell=1}^{\infty}\frac{1}{\left[(\ell-r)^{2}+1\right]^{2}}\right)}$ (71)
$=\frac{\cos(2\pi r)(4\pi\cosh(2\pi)-2\sinh(2\pi))-4\pi+\sinh(4\pi)}{2(\cos(2\pi r)-\cosh(2\pi))^{2}}\,.$ (72)
We can numerically compare the two methods to see these differences in error, as shown in Fig. 5.
V Connection to Scattering Theory
If the system Hamiltonian in (1) is non-interacting, we can fully describe the steady-state transport properties by specifying the transmission probability for an electron with energy $\omega$ to tunnel from bath $\beta$ to bath $\alpha$, which is given by Blanter and Büttiker (2000):
$T_{\alpha\beta}(\omega)=\operatorname{Tr}\big\{J_{\alpha}(\omega)G(\omega)J_{\beta}(\omega)G^{\dagger}(\omega)\big\}\,.$ (73)
Here $\alpha,\beta\in\{L,R\}$, so this includes both transmission and reflection. The quantity $G$ is the single particle delayed non-equilibrium Green’s function (NEGF), which is an $n_{S}\times n_{S}$ matrix given by
$G(\omega)=\frac{1}{\omega-H_{S}-\Sigma(\omega)}\,$, (74)
where $\Sigma(\omega)=\sum_{\alpha}\Sigma_{\alpha}(\omega)$ is the self-energy
$\Sigma_{\alpha}(\omega)=-\frac{iJ_{\alpha}(\omega)}{2}+\Upsilon_{\alpha}(\omega)\,,\quad\Upsilon_{\alpha}(\omega):=\mathcal{P}\intop\frac{d\Omega}{2\pi}\frac{J_{\alpha}(\Omega)}{\omega-\Omega}\,.$ (75)
The imaginary part is the spectral density and the real part $\Upsilon_{\alpha}$ is the level shift.
Under the pseudomode mapping with Hamiltonian (I.2), we treat the system and the two set of pseudomodes as an extended system. We can then study the transmission probability for an electron with energy $\omega$ to tunnel between residual bath $(\alpha,k)$ and residual bath $(\beta,q)$. This therefore adds an additional layer of resolution to the tunneling probabilities. To do that, we first recall that the effects of the residual baths in each pseudomode are captured by the spectral densities:
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p}\tau_{\alpha k,p}^{2}\delta(\omega-\omega_{p})\,$, (76)
as well as the corresponding level shifts $\Upsilon_{\alpha k}^{\rm res}$, defined as in (75). Here, because each pseudomode is coupled to its own reservoir, (76) is a scalar, rather than a matrix like (2).
We can now use a formula akin to (73) to write the transmission probability $T_{\alpha k,\beta q}^{\rm res}$ between residual baths. To do that, introduce the NEGF of the extended system:
$\mathcal{G}(\omega)=\frac{1}{\omega-Q(\omega)}\,$, (77)
where
$Q(\omega)=\begin{bmatrix}\Lambda_{L}+\Sigma_{L}(\omega)&0&\zeta_{L}\\[5.69046pt] 0&\Lambda_{R}+\Sigma_{R}(\omega)&\zeta_{R}\\[5.69046pt] \zeta_{L}^{\dagger}&\zeta_{R}^{\dagger}&H_{S}\end{bmatrix}\,$, (78)
which is of size $n_{L}+n_{R}+n_{S}$. The matrices $\zeta_{\alpha}$ are of size $n_{\alpha}\times n_{S}$. The matrices $\Lambda_{\alpha}$ are the same ones that appear in (I.2). And the matrices $\Sigma_{\alpha}$ are diagonal, with entries $iJ_{\alpha k}^{\rm res}/2+\Upsilon_{\alpha k}^{\rm res}$ in each diagonal.
The transmission probabilities will then have the form
$T_{\alpha k,\beta q}^{\rm res}(\omega)=\operatorname{Tr}\big\{\mathcal{X}_{\alpha k}(\omega)\mathcal{G}(\omega)\mathcal{X}_{\beta q}(\omega)\mathcal{G}^{\dagger}(\omega)\big\}\,.$ (79)
Once again, this describes transmissions from left to right, as well as reflections. Moreover, it also describes reflections to the same pseudomode, as well as to the different pseudomodes belonging to the same bath. The matrices $\mathcal{X}_{\alpha k}$ in (79) are diagonal, of size $d_{L}+d_{R}+n_{S}$, with a single non-zero entry $J_{\alpha k}^{\rm res}$ at position $(\alpha,k)$ (in the same block structure as (78)).
Our next goal is to express (79) in terms of matrices with dimension $n_{S}$ of the system alone. This will allow us to compare the results with those of (73). Consider the Green’s function (77) and parametrize it in blocks as
$\mathcal{G}=\begin{bmatrix}\mathcal{G}_{LL}&\mathcal{G}_{LR}&\mathcal{G}_{LS}\\ \mathcal{G}_{RL}&\mathcal{G}_{RR}&\mathcal{G}_{RS}\\ \mathcal{G}_{SL}&\mathcal{G}_{SR}&\mathcal{G}_{S}\end{bmatrix}\,$. (80)
Using the fact that the matrices $\mathcal{X}_{\alpha k}$ in (79) are all sparse, we can write down four cases of interest:
$T_{Lk,Lq}^{\rm res}$ $=~\operatorname{Tr}\big\{\hat{J}_{Lk}^{\rm res}\mathcal{G}_{LL}\hat{J}_{Lq}^{\rm res}\mathcal{G}_{LL}^{\dagger}\big\}\,$, (81)
$T_{Rk,Rq}^{\rm res}$ $=~\operatorname{Tr}\big\{\hat{J}_{Rk}^{\rm res}\mathcal{G}_{RR}\hat{J}_{Rq}^{\rm res}\mathcal{G}_{RR}^{\dagger}\big\}\,$, (82)
$T_{Lk,Rq}^{\rm res}$ $=~\operatorname{Tr}\big\{\hat{J}_{Lk}^{\rm res}\mathcal{G}_{LR}\hat{J}_{Rq}^{\rm res}\mathcal{G}_{LR}^{\dagger}\big\}\,$, (83)
$T_{Rk,Lq}^{\rm res}$ $=~\operatorname{Tr}\big\{\hat{J}_{Rk}^{\rm res}\mathcal{G}_{RL}\hat{J}_{Lq}^{\rm res}\mathcal{G}_{RL}^{\dagger}\big\}\,$, (84)
where the dependence on $\omega$ has been omitted for clarity. Here $\hat{J}_{\alpha k}^{\rm res}$ is a matrix with dimension $n_{\alpha}\times n_{\alpha}$ having only one entry $J_{\alpha k}^{\rm res}$ in position $k$ of the diagonal. These results can also be written more compactly as
$T_{\alpha k,\beta q}^{\rm res}=\operatorname{Tr}\big\{\hat{J}_{\alpha k}^{\rm res}\mathcal{G}_{\alpha\beta}\hat{J}_{\beta q}^{\rm res}\mathcal{G}_{\alpha\beta}^{\dagger}\big\}\,.$ (85)
To proceed, we therefore need the blocks $\mathcal{G}_{LL},\mathcal{G}_{RR},\mathcal{G}_{LR},\mathcal{G}_{RL}$, which we will express in terms of $\mathcal{G}_{S}$, the effective system NEGF (the (3,3) entry in (80)).
(77) means that $\mathcal{G}$ is the solution of the equations
$(\omega-Q)\mathcal{G}$ $=1,$ (86)
$\mathcal{G}(\omega-Q)$ $=1.$ (87)
Using the block forms in (80) and (78) this yields the following system of coupled equations. First, for the (1,1) and (2,2) diagonal entries we get
$(\omega-\Lambda_{\alpha}-\Sigma_{\alpha})\mathcal{G}_{\alpha\alpha}$ $=1+\zeta_{\alpha}\mathcal{G}_{S\alpha}\,$, (88)
$\mathcal{G}_{\alpha\alpha}(\omega-\Lambda_{\alpha}-\Sigma_{\alpha})$ $=1+\mathcal{G}_{\alpha S}\zeta_{\alpha}^{\dagger}\,.$ (89)
Next, for the (3,3) entries:
$(\omega-H_{S})\mathcal{G}_{S}$ $=1+\big\{\zeta_{L}^{\dagger}\mathcal{G}_{LS}+\zeta_{R}^{\dagger}\mathcal{G}_{RS}\big\}\,$, (90)
$\mathcal{G}_{S}(\omega-H_{S})$ $=1+\big\{\mathcal{G}_{SL}\zeta_{L}+\mathcal{G}_{SR}\zeta_{R}\big\}\,.$ (91)
Finally, for the off-diagonal entries we get:
$(1,2):\quad\mathcal{G}_{LR}$ $=\mathcal{F}_{L}\zeta_{L}\mathcal{G}_{SR}=\mathcal{G}_{LS}\zeta_{R}^{\dagger}\mathcal{F}_{L}$ (92)
$(2,1):\quad\mathcal{G}_{RL}$ $=\mathcal{F}_{L}\zeta_{R}\mathcal{G}_{SL}=\mathcal{G}_{RS}\zeta_{L}^{\dagger}\mathcal{F}_{L}$ (93)
$(1,3):\quad\mathcal{G}_{LS}$ $=\mathcal{F}_{L}\zeta_{L}\mathcal{G}_{S}=\big\{\mathcal{G}_{LL}\zeta_{L}+\mathcal{G}_{LR}\zeta_{R}\big\}\frac{1}{\omega-H_{S}}$ (94)
$(3,1):\quad\mathcal{G}_{SL}$ $=\frac{1}{\omega-H_{S}}\big\{\zeta_{L}^{\dagger}\mathcal{G}_{LL}+\zeta_{R}^{\dagger}\mathcal{G}_{RL}\big\}=\mathcal{G}_{S}\zeta_{L}^{\dagger}\mathcal{F}_{L}$ (95)
$(2,3):\quad\mathcal{G}_{RS}$ $=\mathcal{F}_{L}\zeta_{R}\mathcal{G}_{S}=\big\{\mathcal{G}_{RL}\zeta_{L}+\mathcal{G}_{RR}\zeta_{R}\big\}\frac{1}{\omega-H_{S}}$ (96)
$(3,2):\quad\mathcal{G}_{SR}$ $=\frac{1}{\omega-H_{S}}\big\{\zeta_{L}^{\dagger}\mathcal{G}_{LR}+\zeta_{R}^{\dagger}\mathcal{G}_{RR}\big\}=\mathcal{G}_{S}\zeta_{R}^{\dagger}\mathcal{F}_{L}$ (97)
where we defined
$\mathcal{F}_{\alpha}=\frac{1}{\omega-\Lambda_{\alpha}-\Sigma_{\alpha}}\,.$ (98)
Plug (94) and (96) into (90) to find
$\big(\omega-H_{S}-\zeta_{L}^{\dagger}\mathcal{F}_{L}\zeta_{L}-\zeta_{R}^{\dagger}\mathcal{F}_{L}\zeta_{R}\big)\mathcal{G}_{S}=1\,.$ (99)
Hence, the effective system NEGF is given by
$\mathcal{G}_{S}=\frac{1}{\omega-H_{S}-\sum_{\alpha}\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}(\omega)\zeta_{\alpha}}\,.$ (100)
Next we express the four quantities $\mathcal{G}_{LL},\mathcal{G}_{RR},\mathcal{G}_{LR},\mathcal{G}_{RL}$ [which appear in (81)-(84)] in terms of $\mathcal{G}_{S}$. Combining (92) with (97) we get
$\mathcal{G}_{LR}=\mathcal{F}_{L}\zeta_{L}\mathcal{G}_{S}\zeta_{R}^{\dagger}\mathcal{F}_{L}\,.$ (101)
Combining (93) with (95) we get
$\mathcal{G}_{RL}=\mathcal{F}_{L}\zeta_{R}\mathcal{G}_{S}\zeta_{L}^{\dagger}\mathcal{F}_{L}\,.$ (102)
Finally, combining (88) with (95) or (97) we get
$\mathcal{G}_{\alpha\alpha}=\mathcal{F}_{\alpha}+\mathcal{F}_{\alpha}\zeta_{\alpha}\mathcal{G}_{S}\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}\,.$ (103)
In short, we conclude that
$\mathcal{G}_{\alpha\beta}=\delta_{\alpha\beta}\mathcal{F}_{\alpha}+\mathcal{F}_{\alpha}\zeta_{\alpha}\mathcal{G}_{S}\zeta_{\beta}^{\dagger}\mathcal{F}_{\beta}\,$, (104)
which is the general formula connecting all relevant NEGFs to the system NEGF.
As the last step, we plug (104) into (85). First, for $\beta\neq\alpha$ we get
$T_{\alpha k,\beta q}^{\rm res}$ $=\operatorname{Tr}\big\{\hat{J}_{\alpha k}^{\rm res}(\mathcal{F}_{\alpha}\zeta_{\alpha}\mathcal{G}_{S}\zeta_{\beta}^{\dagger}\mathcal{F}_{\beta})\hat{J}_{\beta q}^{\rm res}(\mathcal{F}_{\beta}^{\dagger}\zeta_{\beta}\mathcal{G}_{S}^{\dagger}\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}^{\dagger})\big\}$ (105)
$=\operatorname{Tr}\big\{\big(\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}^{\dagger}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}\zeta_{\alpha})\mathcal{G}_{S}(\zeta_{\beta}^{\dagger}\mathcal{F}_{\beta}\hat{J}_{\beta q}^{\rm res}\mathcal{F}_{\beta}^{\dagger}\zeta_{\beta})\mathcal{G}_{S}^{\dagger}\big\}\,.$
We therefore see that we can write this in a form similar to (73), provided we define $n_{S}\times n_{S}$ dimensional matrices
$\mathcal{J}_{\alpha k}^{+}=\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}^{\dagger}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}\zeta_{\alpha}\,,\qquad\mathcal{J}_{\alpha k}^{-}=\zeta_{\alpha}^{\dagger}\mathcal{F}_{\alpha}\hat{J}_{\alpha k}^{\rm res}\mathcal{F}_{\alpha}^{\dagger}\zeta_{\alpha}\,.$ (106)
Then
$T_{\alpha k,\beta q}^{\rm res}=\operatorname{Tr}\big\{\mathcal{J}_{\alpha k}^{+}\mathcal{G}_{S}\mathcal{J}_{\beta q}^{-}\mathcal{G}_{S}^{\dagger}\big\}\,.$ (107)
As usual, we now take the pseudomodes’ residual spectral densities to be constant:
$J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha k}\,.$ (108)
Then, if we let $\mathcal{J}_{\alpha}^{+}(\omega)=\sum_{k}\mathcal{J}_{\alpha k}(\omega)$, we can see that, starting from (106):
$\mathcal{J}^{\pm}_{\alpha}$ $=\zeta_{\alpha}^{\dagger}\frac{1}{\omega+iW^{\dagger}_{\alpha}}\Gamma_{\alpha}\frac{1}{\omega-iW_{\alpha}}\zeta_{\alpha}$ (109)
$=\zeta_{\alpha}^{\dagger}\frac{1}{\omega+iW_{\alpha}^{\dagger}}(W_{\alpha}+W_{\alpha}^{\dagger})\frac{1}{\omega-iW_{\alpha}}\zeta_{\alpha}$ (110)
$=i\zeta^{\dagger}\left(\frac{1}{iW_{\alpha}-\omega}+\frac{1}{iW_{\alpha}^{\dagger}+\omega}\right)\zeta_{\alpha}$ (111)
$=2\mathfrak{Im}\left(\zeta_{\alpha}^{\dagger}\frac{1}{iW_{\alpha}-\omega}\zeta_{\alpha}\right)\,.$ (112)
So as per (16) we have that
$\sum_{k}\left[\mathcal{J}_{\alpha k}^{\pm}(\omega)\right]_{ij}=J^{\rm eff}_{\alpha,ij}(\omega)\,.$ (113)
Likewise, if we have constant residual spectral densities (108) then $\Sigma_{\alpha}=i\Gamma_{\alpha}/2$}, and so the effective system Green’s function is:
$\mathcal{G}_{S}(\omega)=\frac{1}{\omega-H_{S}-\sum_{\alpha}\zeta_{\alpha}^{\dagger}\frac{1}{\omega+iW_{\alpha}}\zeta}\,.$ (114)
Recall that the true Green’s function is:
$G(\omega)=\frac{1}{\omega-H_{S}+\sum_{\alpha}\frac{iJ_{\alpha}(\omega)}{2}-\mathcal{P}\intop\frac{d\Omega}{2\pi}\frac{J_{\alpha}(\Omega)}{\omega-\Omega}}\,.$ (115)
The principal value $\Upsilon_{\alpha}$ in the NEGF arises from forcing the memory kernel to be causal:
$\intop dt\chi_{\alpha}(t)e^{i\omega t}\theta(t)=\frac{J_{\alpha}(\omega)}{2}-\mathcal{P}\intop\frac{d\Omega}{2\pi}\frac{J_{\alpha}(\Omega)}{\omega-\Omega}\,.$ (116)
where $\theta(t)$ is the Heaviside step function. We can see the equivalence between (115) and (116) by noticing that
$\intop dt\chi_{\alpha,ij}^{\rm eff}(t)e^{i\omega t}\theta(t)=\intop dt\zeta_{\alpha}^{\dagger}e^{(W_{\alpha}+i\omega)t}\zeta_{\alpha}=\zeta_{\alpha}\frac{1}{W_{\alpha}+i\Omega}\zeta_{\alpha}\,.$ (117)
We can then write in the pseudomode picture:
$T^{\rm eff}_{\alpha\beta}(\omega)=\operatorname{Tr}\left\{J^{\rm eff}_{\alpha}(\omega)\mathcal{G}_{S}(\omega)J^{\rm eff}_{\beta}(\omega)\mathcal{G}_{S}^{\dagger}(\omega)\right\}\,$, (118)
$\mathcal{G}_{S}(\omega)=\frac{1}{\omega-H_{S}+\sum_{\alpha}\frac{iJ_{\alpha}^{\rm eff}(\omega)}{2}-\mathcal{P}\intop\frac{d\Omega}{2\pi}\frac{J^{\rm eff}_{\alpha}(\Omega)}{\omega-\Omega}}\,$, (119)
where $\mathcal{G}_{S}$ is just the NEGF that we would calculate from the system interacting with a set of baths with spectral densities $J^{\rm eff}_{\alpha}$. Thus, we can see that so long as the spectral densities (or correlation functions) match exactly, the transmission functions will also match exactly.
VI Conclusion
The method of pseudomodes is a useful approach to studying the dynamics of open quantum systems with strong coupling or structured environments. We have explored a number of subtleties that arise in the implementation of this technique, including the differences in effective spectral densities for coupled and uncoupled pseudomodes, and the appearance of new terms in the effective spectral density when the pseudomodes’ non-Hermitian Hamiltonian is non-diagonalizable. We have explored the pseudomode inversion problem and presented for the first time a method for finding pseudomode parameters that exactly match an effective fit to a spectral density. Our inversion method also sheds further light on the vast number of free parameters involved in the pseudomode method. We have also shown a connection between pseudomodes and scattering theory for non-interacting systems and shown that they produce the same effective spectral densities.
While the method we present for inversion is exact, it is not guaranteed to provide a pseudomode configuration with positive damping rates. We have explored the constraints on obtaining positive rates in the limited case of 2 pseudomodes, and shown that it is possible in that case to obtain positive rates if and only if the effective spectral density is positive everywhere. However, it is currently unclear whether this result will always hold in general.
Acknowledgments
This research is primarily supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES) under Award No. DE-SC0025516. L.P.B. acknowledges Research Ireland for support through the Frontiers for the Future project. Optimization was performed with the Optim.jl Julia package Mogensen and Riseth (2018).
Appendix A Derivation of the Effective Spectral Density
In this appendix we derive the forms of the effective spectral density and effective memory kernel given in (16) and (14), respectively, following Ref. Ferracin et al. (2024). Consider the system described in section I.2 with Hamiltonian given by (I.2) in the special case where the residual baths have constant spectral densities $J^{\rm res}_{\alpha k}(\omega)=\Gamma_{\alpha k}$ and the Fermi functions of the original (physical) baths are $f_{\alpha}$. We then model the residual baths using a local master equation of the GKSL form. The Hamiltonian is now:
$H=H_{S}(t)+\sum_{\alpha\in\{L,R\}}\sum_{k,q}\Lambda_{\alpha,kq}a_{\alpha k}^{\dagger}a_{\alpha q}+$ (120)
$\sum_{\alpha\in\{L,R\}}\sum_{i,k}\big(\zeta_{\alpha i,k}a_{\alpha k}^{\dagger}c_{i}+\zeta_{\alpha i,k}^{*}c_{i}^{\dagger}a_{\alpha k}\big)\,$,
such that the total time evolution for an operator $\mathcal{O}$ in the evolving-operators picture is:
$\frac{d\mathcal{O}}{dt}$ $=\mathcal{L}^{\prime}\mathcal{O}=i[H,\mathcal{O}]+$ (121)
$\sum_{\alpha\in\{L,R\}}\sum_{k}\Gamma_{\alpha k}\left(f_{\alpha k}D^{\prime}[a_{\alpha k}^{\dagger}]\mathcal{O}+(1-f_{\alpha k})D^{\prime}[a_{\alpha k}]\mathcal{O}\right)\,$,
where the adjoint dissipator $D^{\prime}[L]\mathcal{O}=L^{\dagger}\mathcal{O}L-\frac{1}{2}\{L^{\dagger}L,\mathcal{O}\}$.
Since the spectral density and memory kernel are independent of reservoir temperature or chemical potential, we perform the derivation for the special case of a competely empty reservoir: one whose Fermi function $f_{\alpha}(\omega)=0$. In this special case, the bath correlation function $C_{\alpha ij}^{+}(t)$ is zero at all times, and $C_{\alpha ij}^{-}(t)$ is identical to the memory kernel.
$\chi_{\alpha,ij}(t)\big|_{f(\omega)=0}=C_{\alpha,ij}^{-}(t)=\left\langle B_{\alpha i}(t)B_{\alpha j}^{\dagger}(0)\right\rangle\,.$ (122)
Additionally, for simplicity, since each bath will have its own independent correlation functions, we consider the case of only one reservoir and so can suppress the index $\alpha$. We can then evaluate the memory kernel matrix $\chi(t)$ by way of the expectation value in the evolving-operators picture (10), where $|{\mathbf{0}}\rangle$} is the full vacuum state of the bath:
$\chi_{ij}(t)$ $=\left\langle B_{i}(t)B_{j}^{\dagger}(0)\right\rangle$ (123)
$=\sum_{k}\sum_{l}\zeta_{i,k}\zeta_{j,l}^{*}{\langle{\mathbf{0}}|}a_{k}(t)a_{l}^{\dagger}(0){|{\mathbf{0}}\rangle}$ (124)
$=\sum_{k}\sum_{l}\zeta_{i,k}\zeta_{i,l}^{*}{\langle{\mathbf{0}}|}a_{k}(t){|{\mathbf{1}_{l}}\rangle}\,$, (125)
where ${|{\mathbf{1}_{l}}\rangle}=a_{l}{|{\mathbf{0}}\rangle}$. Now define a vector $v_{i}$ such that its components are $v_{i,j}(t)={\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{j}}\rangle}$ and consider:
$\frac{dv_{i,j}}{dt}$ $={\langle{0}|}\dot{a}_{i}(t){|{1_{j}}\rangle}={\langle{0}|}\mathcal{L}^{\prime}({a}_{i}(t)){|{1_{j}}\rangle}\,$, (126)
where $\mathcal{L}^{\prime}$ is the adjoint Liouvillian that arises from evolution under (121). The two terms are then, where $\mathcal{D}^{\prime}$ represents only the dissipator part of (121):
${\langle{\mathbf{0}}|}[H_{E},a_{i}(t)]{|{\mathbf{1}_{j}}\rangle}$ $=0+\sum_{m=1}^{n_{\alpha}}\Lambda_{mj}{\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{m}}\rangle}\,$, (127)
${\langle{\mathbf{0}}|}\mathcal{D}^{\prime}(a_{i}(t)){|{1_{j}}\rangle}$ $=-\frac{1}{2}\sum_{m=1}^{n_{\alpha}}\Gamma_{mj}{\langle{\mathbf{0}}|}a_{i}(t){|{\mathbf{1}_{m}}\rangle}\,.$ (128)
So combining these and putting them into vector form, we have:
$\frac{dv_{i}}{dt}=\left(-i\Lambda-\frac{1}{2}\Gamma\right)^{T}v^{i}(t)\,.$ (129)
If we let $W_{\alpha}=\left(-i\Lambda_{\alpha}-\frac{1}{2}\Gamma_{\alpha}\right)$, noting that $iW$ is the non-Hermitian Hamiltonian for this model, then this equation has the solution:
$v_{i,j}(t)=\left(e^{tW}\right)_{ij}\,.$ (130)
Finally, the memory kernel is then given by the matrix equation:
$\chi_{ij}(t)=\zeta_{i}^{\dagger}e^{Wt}\zeta_{j}\,$, (131)
as given by (14). Taking the Fourier transform yields the result for the spectral density given in (16).
$J_{\alpha ij}(\omega)=2\mathfrak{Im}\left(\zeta_{\alpha i}^{\dagger}\frac{1}{iW-\omega}\zeta_{\alpha j}\right)\,.$ (132)
Appendix B Inversion Problem for 2 Pseudomodes
In this appendix we explore the method presented in section III for the simple case of two pseudomodes. Suppose we obtain a 2-mode fit from Prony’s method with eigenvalues $\varepsilon_{1},\varepsilon_{2},\gamma_{1},\gamma_{2}$ and amplitudes $\kappa_{1},\kappa_{2}$. First note that to satisfy (25), the $\kappa$’s must be of the following form:
$\kappa_{1}=\alpha_{1}+i\beta\,,\quad\kappa_{2}=\alpha_{2}-i\beta\,$, (133)
where $\alpha_{1},\alpha_{2},\beta$ are real numbers and $\alpha_{1}+\alpha_{2}>0$. In (45), we fix
$u_{j}=1\,,\quad v_{j}=\kappa_{j}^{\star}$ (134)
Since the matrices here are only 2x2, we can then construct $S^{\dagger}S$ directly from (46) as:
$S^{\dagger}S$ $=\begin{bmatrix}\kappa_{1}^{*}/2&\kappa_{1}^{*}/2\\ \kappa_{2}^{*}/2&\kappa_{2}^{*}/2\end{bmatrix}+\begin{bmatrix}a&-a\\ b&-b\end{bmatrix}\,$. (135)
Enforcing positivity then requires:
$(S^{\dagger}S)_{11}>0$ $\implies a=\frac{\kappa_{1}}{2}+a^{\prime}\,,\,a^{\prime}>-\alpha_{1}$ (136)
$(S^{\dagger}S)_{22}>0$ $\implies b=-\frac{\kappa_{2}}{2}+b^{\prime}\,,\,b^{\prime}>-\alpha_{2}\,.$ (137)
and then for Hermiticity we must have $b^{\prime}=a^{\prime}$. We thus obtain $S^{\dagger}S$ in terms of one free parameter $a^{\prime}>\max\{0,-\alpha_{1},-\alpha_{2}\}$ as:
$S^{\dagger}S=\begin{bmatrix}\alpha_{1}+a^{\prime}&i\beta+a^{\prime}\\ -i\beta-a^{\prime}&\alpha_{2}+a^{\prime}\end{bmatrix}\,$. (138)
The final constraint to enforce $S^{\dagger}S>0$ is that we must have
$\det(S^{\dagger}S)=a^{\prime}(\alpha_{1}+\alpha_{2})+\alpha_{1}\alpha_{2}-\beta^{2}>0\,$, (139)
and since we must have $\alpha_{1}+\alpha_{2}>0$ from (25) we can always choose $a^{\prime}$ large enough that (139) is satisfied. Since $S^{\dagger}S$ is $2\times 2$ we can express $S=\sqrt{S^{\dagger}S}$ analytically as Levinger (1980):
$S$ $=\frac{1}{\sqrt{2a+\alpha_{1}+\alpha_{2}+2\sqrt{\det(S^{\dagger}S)}}}\quad\times$ (140)
$\begin{bmatrix}a^{\prime}+\alpha_{1}+\sqrt{\det(S^{\dagger}S)}&-a^{\prime}-i\beta\\ -a^{\prime}+i\beta&a^{\prime}+\alpha_{2}+\sqrt{\det(S^{\dagger}S)}\end{bmatrix}\,$.
Using (52) we can find $\tilde{\Gamma}$ and compute its eigenvalues, which will be the same as the eigenvalues of the diagonal $\Gamma$ obtained in (56). Those eigenvalues have the form:
$\frac{\gamma_{1}+\gamma_{2}}{2}\pm\frac{\sqrt{c_{1}(\gamma_{1}-\gamma_{2})^{2}+c_{2}(\varepsilon_{1}-\varepsilon_{2})^{2}}}{c_{3}}\,$, (141)
where $c_{1}$, $c_{2}$, $c_{3}$ are all positive and depend only on $\alpha_{1}$, $\alpha_{2}$, $\beta$, and $a^{\prime}$. From this equation we can see that any choice of $a^{\prime}$ that makes one of the rates smaller will make the other larger. We can also see that in order to have both rates be positive, the second term must be smaller than the first. Thus, given some $\alpha_{1},\alpha_{2},\beta$ from a 2-mode Prony’s method fit we can immediately tell whether it will be possible for that fit to have both rates be positive by looking at (141) and asking whether there is a choice of $a^{\prime}$ such that the second term is smaller than the first.
We can simplify further in the case that we have a symmetric spectral density, such that $J(\omega)=J(-\omega)$. In that case we expect a fit from Prony’s method to follow the same symmetry, i.e,
$\gamma_{1}=\gamma_{2}\,,\quad\varepsilon_{1}=-\varepsilon_{2}\,,\quad\alpha_{1}=\alpha_{2}\,.$ (142)
In this case, (141) simplifies greatly to (removing the redundant subscripts)
$\gamma\pm\frac{2\sqrt{a^{\prime 2}+\beta^{2}}|\varepsilon|}{\sqrt{2a^{\prime}\alpha+\alpha^{2}-\beta^{2}}}\,.$ (143)
To see under what conditions we can choose $a^{\prime}$ such that both rates are positive, we can then simply solve
$\frac{2\sqrt{a^{\prime 2}+\beta^{2}}\left|\varepsilon\right|}{\sqrt{2a^{\prime}\alpha+\alpha^{2}-\beta^{2}}}=r\gamma_{1}\,$, (144)
for any $0\leq r\leq 1$ and see under what conditions we find a real solution for $a^{\prime}$. Choosing $r=1$ makes one of the rates zero and maximizes the other, while $r=0$ would make both rates identically $\gamma_{1}$. Squaring both sides of (144), we can rearrange to obtain an equation that is quadratic in $a^{\prime}$; the solutions are then
$a^{\prime}=\frac{\alpha r^{2}\gamma_{1}^{2}\pm\sqrt{\left(r^{2}\gamma^{2}+4\varepsilon^{2}\right)\left(\alpha^{2}r^{2}\gamma^{2}-4\beta^{2}\varepsilon^{2}\right)}}{4\varepsilon^{2}}\,$, (145)
and there is only a real solution if
$\alpha^{2}r^{2}\gamma^{2}>4\beta^{2}\varepsilon^{2}\,.$ (146)
This condition is least restrictive when $r=1$ and so we can see that there is only a choice of $a^{\prime}$ that will result in a $\Gamma$ matrix that is positive if the fit parameters satisfy
$\frac{\alpha^{2}}{\beta^{2}}>4\frac{\varepsilon^{2}}{\gamma^{2}}\,.$ (147)
We now compare this bound with the bounds on fit parameters such that the resulting effective spectral density is positive everywhere. The effective spectral density (26) corresponding to a fit satisfying (142) is
$J_{\rm eff}(\omega)=\frac{\alpha\gamma+2\beta(\omega-\varepsilon)}{(\omega-\varepsilon)^{2}+(\gamma/2)^{2}}+\frac{\alpha\gamma-2\beta(\omega+\varepsilon)}{(\omega+\varepsilon)^{2}+(\gamma/2)^{2}}\,.$ (148)
We assume without loss of generality that $\varepsilon>0$; we then might have $\beta$ to be positive or negative. Computing $dJ_{\rm eff}(\omega)/d\omega=0$, we find up to five potential critical points. One of them is at $\omega=0$, and the others are:
$\omega_{c,\pm,\pm^{\prime}}=\pm\frac{1}{2}\,\times$ (149)
$\sqrt{\frac{\pm^{\prime}4\gamma|\varepsilon|\sqrt{\left(\alpha^{2}+\beta^{2}\right)\left(\gamma^{2}+4\varepsilon^{2}\right)}-\left(\gamma^{2}+4\varepsilon^{2}\right)(\alpha\gamma-2\beta\varepsilon)}{\alpha\gamma+2\beta\varepsilon}}\,.$
Starting with the critical point at zero, we see
$J_{\rm eff}(0)=\frac{2\alpha\gamma-4\beta\varepsilon}{\varepsilon^{2}+\gamma^{2}/4}\,$, (150)
and so for this to be positive we require
$\alpha\gamma>2\beta\varepsilon\,.$ (151)
Which of the other possible critical points are real will depend on the sign of $\alpha\gamma+2\beta\varepsilon$. We first consider the case that it is positive, in which case there can only possibly be critical points for $\pm^{\prime}=+$. We then see that
$J_{\rm eff}(\omega_{c,+,+})=\frac{2\alpha\varepsilon-\beta\gamma+\sqrt{(\alpha^{2}+\beta^{2})(\gamma^{2}+4\varepsilon^{2})}}{\gamma\varepsilon}\,$, (152)
which is always positive, since $\sqrt{(\alpha^{2}+\beta^{2})(\gamma^{2}+4\varepsilon^{2})}>|\beta|\gamma$. Thus, $J_{\rm eff}$ will always be positive at these points (if they are real).
Finally, we consider the case that $\alpha\gamma+2\beta\varepsilon<0$, in which case we know $\beta<0$ and there will be at least two critical points where $\pm^{\prime}=$-. We then obtain
$J_{\rm eff}(\omega_{c,+,-})=\frac{2\alpha\varepsilon-\beta\gamma-\sqrt{(\alpha^{2}+\beta^{2})(\gamma^{2}+4\varepsilon^{2})}}{\gamma\varepsilon}\,$, (153)
and for this to be positive, we have:
$2\alpha\varepsilon+|\beta|\gamma$ $>\sqrt{(\alpha^{2}+\beta^{2})(\gamma^{2}+4\varepsilon^{2})}\,.$ (154)
Squaring both sides and rearranging, we find we must have
$(\alpha\gamma-2\beta\varepsilon)^{2}<0\,$, (155)
which is impossible. So we have the additional constraint that we must have $\alpha\gamma+2\beta\varepsilon>0$. Combining this with (151), we see that we can write
$\alpha\gamma>2|\beta|\varepsilon\,$, (156)
or, to mirror (147):
$\frac{\alpha^{2}}{\beta^{2}}>4\frac{\varepsilon^{2}}{\gamma^{2}}\,.$ (157)
This exactly matches the bound (147)! This means that, given a set of fit parameters $\alpha,\beta,\gamma,\varepsilon$, we can find an $a^{\prime}$ such that $\Gamma>0$ if and only if the effective spectral density from those fit parameters is positive.
Appendix C Non-Diagonalizable Configuration as the Limit of a Diagonalizable One
Consider the 2x2 non-diagonalizable block from section II.3 where we set $\varepsilon=\eta=0$, and apply a small perturbation $\epsilon$ to the coupling between the two modes:
$W=\begin{bmatrix}0&-i\delta(1+\epsilon)\\ -i\delta(1+\epsilon)&-2\delta\end{bmatrix}\,$. (158)
$W$ is diagonalizable so long as $\epsilon\neq 0$, and its eigenvalues are
$\delta\left(1\pm i\sqrt{2\epsilon(1+\epsilon)}\right)\,.$ (159)
We then compute the memory kernel using (14) and obtain:
[…]
$\frac{\left(\zeta_{1}^{2}-\zeta_{2}^{2}-2i\zeta_{2}\zeta_{1}(\epsilon+1)\right)\sin\left(\delta t\sqrt{\epsilon(\epsilon+2)}\right)}{\sqrt{\epsilon(\epsilon+2)}}\Bigg]\,.$
In the limit of $\epsilon\to 0$, the first two terms become $(\zeta_{1}^{2}+\zeta_{2}^{2})e^{-t\delta}$}. The third term is proportional to $\frac{\sin(x)}{x}$ and so in the limit we get $t\delta e^{-t\delta}(\zeta_{1}-i\zeta_{2})^{2}$, the signature we would expect from a non-diagonalizable $W$. When we compute the resulting spectral density, we obtain:
$J(\omega)=\frac{4\delta\left(\delta\zeta_{1}+\zeta_{2}\omega\right){}^{2}}{\left(\delta^{2}+\omega^{2}\right)^{2}}\,.$ (161)
Note that in the case we take $\zeta_{2}=0$ we recover exactly the squared Lorentzian that we would expect:
$J(\omega)=\frac{4\delta^{3}\zeta_{1}^{2}}{\left(\delta^{2}+\omega^{2}\right)^{2}}\,.$ (162)
If instead we construct the spectral density before taking the limit, then we can define $\kappa_{k}=\sum_{i=1}^{2}\sum_{j=1}^{2}\zeta_{i}^{*}\zeta_{j}S_{ik}S^{-1}_{kj}=a_{k}+ib_{k}$ as in section II.2, where $S$ is the matrix that diagonalizes $W$. We obtain:
$a_{1}$ $=\frac{1}{2}\left(\zeta_{1}^{2}+\zeta_{2}^{2}+\frac{2\zeta_{1}\zeta_{2}(\epsilon+1)}{\sqrt{\epsilon(\epsilon+2)}}\right)\,$, (163)
$a_{2}$ $=\frac{1}{2}\left(\zeta_{1}^{2}+\zeta_{2}^{2}-\frac{2\zeta_{2}\zeta_{1}(\epsilon+1)}{\sqrt{\epsilon(\epsilon+2)}}\right)\,$, (164)
$b_{1}$ $=\frac{\zeta_{1}^{2}-\zeta_{2}^{2}}{2\sqrt{\epsilon(\epsilon+2)}}\,$, (165)
$b_{2}$ $=\frac{\zeta_{2}^{2}-\zeta_{1}^{2}}{2\sqrt{\epsilon(\epsilon+2)}}\,.$ (166)
Note that each of these diverges as $\epsilon\to 0$. However, if we construct the spectral density by
$J(\omega)=\sum_{i=1}^{n_{\alpha}}a_{i}\frac{\gamma_{i}}{(\omega-\varepsilon_{i})^{2}+(\gamma_{i}/2)^{2}}-b_{i}\frac{2(\omega-\varepsilon_{i})}{(\omega-\varepsilon_{i})^{2}+(\gamma_{i}/2)^{2}}\,$, (167)
where $\epsilon_{i},\gamma_{i}$ come from the eigenvalues of $W$, then we obtain:
$J(\omega)=\frac{4\delta\left(\delta\zeta_{1}(\epsilon+1)+\zeta_{2}\omega\right){}^{2}}{\delta^{4}(\epsilon+1)^{4}-2\delta^{2}\omega^{2}(\epsilon(\epsilon+2)-1)+\omega^{4}}\,.$ (168)
Taking the limit of $\epsilon\to 0$, we obtain:
$J(\omega)=\frac{4\delta\left(\delta\zeta_{1}+\zeta_{2}\omega\right){}^{2}}{\left(\delta^{2}+\omega^{2}\right)^{2}}\,$, (169)
exactly as before.
Appendix D Prony’s Method
Suppose we have a memory kernel $\chi(t)$ that we want to approximate as:
$\chi(t)=\sum_{k=1}^{n}\kappa_{k}e^{-\lambda_{k}t}\,$, (170)
for complex parameters $\kappa_{k},\lambda_{k}$, where we want $n$ as small as possible to get a good fit. Prony’s method Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) consists of the following steps:
  1. 1.
    Choose a time interval $\Delta t$ and cutoff time $t_{c}=2N\Delta t$ to sample the correlation function, where $N\geq 2L+1$.
  2. 2.
    Let $\phi_{j}=C(j\Delta t)$ for $j=0,\dots,2N$ and construct the $(N+1)\times(N+1)$ matrix $\mathbf{H}_{ij}=\phi_{i+j+1}$.
  3. 3.
    Compute the Takagi decomposition Ikramov (2012): find vectors $\mathbf{u}_{m}$ such that $\mathbf{H}\mathbf{u}_{m}=\sigma_{m}\mathbf{u}_{m}^{*}$. This is analogous to finding the eigendecomposition of a Hermitian matrix as $\mathbf{H}{|{m}\rangle}=\sigma_{m}{|{m}\rangle}$, but for a complex-symmetric matrix. The matrix $U$ whose columns are $\mathbf{u}_{m}$ is unitary so we could alternately write $H=U\sigma U^{T}$ where we take the transpose of $U$ rather than the adjoint. We order the $\sigma_{m}$ by descending values of $|\sigma_{m}|$ and the $\mathbf{u}_{m}$ accordingly.
  4. 4.
    Define the polynomial $f(z)=\sum_{n=0}^{N}\left(\mathbf{u}_{L}\right)_{n}z^{n}$ and obtain its roots $w_{0},w_{1},\dots w_{N}$ ordered by increasing modulus.
  5. 5.
    Of these roots, the first $L$ will have $|w_{k}|<1$. Write these roots as $r_{k}e^{i\theta_{k}}$ with $\pi<\theta\leq\pi$ and choose the exponents to be $\lambda_{k}=-\frac{2N}{t_{c}}\left[\log(r_{k})+i\theta_{k}\right]$.
  6. 6.
    Obtain $\kappa_{k}$ by least-squares fitting the equations $\phi_{j}=\sum_{k=1}^{L}\kappa_{k}w_{k}^{j}$ for $j=0,1,\dots 2N$.
Fitting with Prony’s method can be quite sensitive to the choice made in step 1. Thus, it is generally most practical to sample a large number of interval choices and then select the choice that gives the effective spectral density that best fits the true spectral density, or to perform optimization on the endpoints of the time interval used.
Appendix E 3x3 non-diagonalizable Configuration
In general, constructing an $n\times n$ non-diagonalizable matrix that is the sum of an anti-Hermitian matrix a and real diagonal matrix, as $W$ is, is a nontrivial task. However, we can use the 2x2 non-diagonalizable block to construct a 3x3 non-diagonalizable block for 3 pseudomodes arranged in a chain. We begin with the 2x2 non-diagonalizable block and expand it to 3x3, leaving the two additional off-diagonal elements as unknown:
$W=(-i\varepsilon-\eta)\mathds{1}+\delta\begin{bmatrix}0&-ix&0\\ -ix&0&-i\\ 0&-i&-2\end{bmatrix}\,$. (171)
We compute the eigenvalues of this matrix as:
$\lambda_{1}(x)$ $=-\frac{2}{3}-\frac{-1+3x^{2}}{3\left(1-18x^{2}+3\sqrt{3}\sqrt{-x^{2}+11x^{4}+x^{6}}\right)^{1/3}}$ (172)
$+$ $\frac{1}{3}\left(1-18x^{2}+3\sqrt{3}\sqrt{-x^{2}+11x^{4}+x^{6}}\right)^{1/3}\,$,
$\lambda_{2}(x)$ $=-\frac{2}{3}+\frac{(1+i\sqrt{3})(-1+3x^{2})}{6\left(1-18x^{2}+3\sqrt{3}\sqrt{-x^{2}+11x^{4}+x^{6}}\right)^{1/3}}$ (173)
- $\frac{1}{6}(1-i\sqrt{3})\left(1-18x^{2}+3\sqrt{3}\sqrt{-x^{2}+11x^{4}+x^{6}}\right)^{1/3}\,$,
$\lambda_{3}(x)$ $=-\frac{2}{3}+\frac{(1-i\sqrt{3})(-1+3x^{2})}{6\left(1-18x^{2}+3\sqrt{3}\sqrt{-x^{2}+11x^{4}+x^{6}}\right)^{1/3}}$ (174)
- $\frac{1}{6}(1+i\sqrt{3})\left(1-18x^{2}+3\sqrt{3}\sqrt{-x^{2}+11x^{4}+x^{6}}\right)^{1/3}\,.$
Plotting the eigenvalues, one can see a point where $\lambda_{1}(x)$ and $\lambda_{3}(x)$ coincide and that it occurs where the derivative of both eigenvalues with respect to $x$ is undefined. Informed by this plot, we are able to find $x$ where both eigenvalues coincide, and can then verify that the resulting $W$ is non-diagonalizable. In particular, we find:
$x=\sqrt{\frac{1}{2}\left(5\sqrt{5}-11\right)}\,$, (175)
and so we find a candidate 3x3 non-diagonalizable $W$ as:
$W=(-i\varepsilon-\eta)\mathds{1}+\delta\begin{bmatrix}0&\sqrt{\frac{1}{2}\left(5\sqrt{5}-11\right)}&0\\ \sqrt{\frac{1}{2}\left(5\sqrt{5}-11\right)}&0&-i\\ 0&-i&-2\end{bmatrix}\,$. (176)
If we then proceed as in the 2x2 case and allow only the first pseudomode to couple to the system and the last pseudomode to couple to the residual bath (so $\eta=0$), we can then compute the resulting effective memory kernel
$\chi_{\rm eff}(t)$ $=|\zeta|^{2}e^{-i\varepsilon t}\left\{\frac{e^{-\left(\sqrt{5}-1\right)\delta t}\left[\left(7\sqrt{5}-17\right)e^{\frac{1}{2}\left(3\sqrt{5}-5\right)\delta t}+8\sqrt{5}-18\right]}{5\left(3\sqrt{5}-7\right)}\right$. (177)
$\qquad\left.+\frac{\delta te^{\frac{1}{2}\left(\sqrt{5}-3\right)\delta t}}{\sqrt{5}}\right\}\,$,
and effective spectral density
$J_{\rm eff}(\omega)=\frac{8(-11+5\sqrt{5})\delta^{5}\lambda_{0}^{2}}{\left[(7-3\sqrt{5})\delta^{2}+2(\omega-\varepsilon)^{2}\right]^{2}\left[2(3-\sqrt{5})\delta^{2}+(\omega-\varepsilon)^{2}\right]}\,$, (178)
for this pseudomode configuration.
We note that there are only terms like $e^{t}$ and $te^{t}$ in the effective memory kernel, since $W$ as given by (176) has two eigenvalues and eigenvectors, rather than just one, and the defective eigenvalue has an algebraic multiplicity of 2.